### v2

@@ Abstract
This paper aims to provide a series of characterizations of the robust isolated calmness of the Karush-Kuhn-Tucker (KKT) mapping for spectral norm regularized convex optimization problems. By establishing the variational properties of the spectral norm function, we directly prove that the KKT mapping is isolated calm if and only if the strict Robinson constraint qualification (SRCQ) and the second order sufficient condition (SOSC) hold. Furthermore, we obtain the crucial result that the SRCQ for the primal problem and the SOSC for the dual problem are equivalent. The obtained results can derive more equivalent or sufficient conditions of the robust isolated calmness of the KKT mapping, thereby enriching the stability theory of spectral norm regularized optimization problems and enhancing the usability of isolated calmness in algorithm applications.
Keywords: Isolated calmness, Second order sufficient condition, Spectral norm, Strict Robinson constraint qualification, Critical cone
@@ 1 Introduction
Stability analysis theory is very crucial in studying the convergence of algorithms. As one of the important concepts in stability properties, isolate calmness (see Definition 2.1) may guarantee the linear convergence rate of some algorithms, such as the alternating direction method of multipliers [9] and the proximal augmented Lagrangian method [10]. There are many publications on studying the isolated calmness of the KKT mapping (ICKKTM) for optimization problems. For the nonlinear semidefinite programming, Zhang and Zhang [11] show that the ICKKTM can be derived from the SOSC and the SRCQ. Zhang, et al. [12] demonstrate that the SOSC and the SRCQ are both sufficient and necessary for the ICKKTM for the nonlinear second order cone programming problem. It is important to note that Ding, Sun and Zhang [13] study the robust isolated calmness of a large class of cone programming problems. They prove that if and only if both the SOSC and the SRCQ hold, the KKT mapping is robustly isolated calm.
@@
Inspired by the work in [16], given the widespread application of spectral norm regularized convex matrix optimization problems, a natural question is whether the results in [16] can be extended to spectral norm regularized convex optimization problems. We provide an answer in this paper. Due to the special structure of the critical cone of spectral norm function, i.e., the expression of the critical cone is not unique, we provide several equivalent or sufficient conditions of the robust ICKKTM. That is to say, part of the conclusions in [16] are still valid for problem (1.1). Compared to [16], the difference in this paper is that we use the established variational properties of the proximal mapping of spectral norm to provide a direct proof of the ICKKTM for problem (1.1), while [16] provides an indirect proof of the isolated calmness based on [13, Theorem 24] regarding the optimization problems with a $\mathcal{C}^{2}$-cone reducible constraint set. In addition, compared to the SOSC, the SRCQ for primal or dual problem is easier to verify, which can enhance the practicality of isolated calmness in algorithm research.
The organization of the subsequent content is as follows. We provide some notations and preliminaries on variational analysis which will be used in the following text in Section 2. The variational properties of the spectral norm function are studied in Section 3, including the relationship between the critical cones of the spectral norm function and its conjugate function, and the explicit expression of the directional derivative of the proximal mapping of the spectral norm function. The results in Section 3 play a vital role in obtaining the important conclusions of this paper. In Section 4, we prove that the primal SRCQ holds if and only if the dual SOSC holds, and thus establish two equivalent conditions and two sufficient condition for the robust ICKKTM for problem (1.1). We conclusion this paper in Section 5.
Some common symbols and notations for matrices are given below:
@@ 2 Notation and Preliminaries
Moreover, we say that $G$ is robustly isolated calm at $\bar{u}$ for $\bar{v}$ if (2.1) holds and for each $u\in\mathcal{U}$, $G(u)\cap\mathcal{V}\neq\emptyset$.
For any closed convex function $g:\mathcal{X}\rightarrow(-\infty,+\infty]$, we know from [18, Proposition 2.58] that $g$ is directionally epidifferentiable. We use $g^{\downarrow}(x;\cdot)$ to denote the directional epiderivatives of $g$. Further, if $g^{\downarrow}(x;d)$ is finite for $x\in{\rm dom}\,g$ and $d\in\mathcal{X}$, we define the lower second order directional epiderivative of $g$ for any $w\in\mathcal{X}$ by
@@ 3 Variational Analysis of the Spectral Norm Function
In this section, we will provide a direct expression for the directional derivative of the proximal mapping of the spectral norm function. Then discuss the relationship between directional derivatives of the proximal mapping, the so called “sigma term” and the critical cones.
@@
where $U\in\mathcal{O}^{m}$ and $[V_{1}\quad V_{2}]\in\mathcal{O}^{n}$ with $V_{1}\in\Re^{n\times m}$ and $V_{2}\in\Re^{n\times(n-m)}$. Define the following three index sets
@@
where $\lambda^{*}>0$ and satisfies $\sum\limits_{i=1}^{m}[\sigma_{i}(Q)-\lambda^{*}]_{+}=1$ (for any $x\in\Re$, $[x]_{+}$ means the nonnegative part of $x$). Then, we can conclude that $\kappa(\sigma)=(\phi(\sigma_{1}),\phi(\sigma_{2}),\cdots,\phi(\sigma_{m}))$ and
@@
$\alpha:=\bigcup_{k=1}^{\tilde{r}}a_{k},\quad\beta:=\bigcup_{\tilde{r}+1}^{r+1}a_{k}.$ (3.4)
@@
$\left\{\begin{array}[]{l}\Xi_{1}(Q,D):=\Theta^{2}_{\alpha\alpha}(Q)\circ T(D_{1})_{\alpha\alpha},\\[15.0pt] \Xi_{2}(Q,D):=\Theta^{1}_{\alpha\beta}(Q)\circ S(D_{1})_{\alpha\beta}+\Theta^{2}_{\alpha\beta}(Q)\circ T(D_{1})_{\alpha\beta},\\[15.0pt] \Xi_{3}(Q,D):=(\Theta^{1}_{\alpha\beta}(Q))^{T}\circ S(D_{1})_{\beta\alpha}+(\Theta^{2}_{\alpha\beta}(Q))^{T}\circ T(D_{1})_{\beta\alpha},\\[15.0pt] \Xi_{4}(Q,D):=\Theta_{\alpha c}(Q)\circ D_{\alpha c}.\\ \end{array}\right.$ (3.11)
@@
Propositions 10 and Proposition 12 in [25] characterized the critical cones of $\theta$ and $\theta^{*}$ defined in (3.8) and (3.9), respectively. We summarize the results as follows.
@@
    where $\lambda(Z)$ denote the real eigenvalue vector of symmetric matrix $Z$ being arranged in non-increasing order.
@@
    If $H\in\mathcal{C}_{\theta}(P,W)$ and $\psi^{*}_{(P,H)}(W)=0$, then $H\in(\mathcal{C}_{\theta^{*}}(P,W))^{\circ}$.
@@
Then, it follows from (ii) of Proposition 3.1 that $H\in\mathcal{C}_{\theta^{*}}(W,P)$ and $\varphi^{*}_{(W,H)}(P)=0$ are equivalent to
@@
which means that $H\in(\mathcal{C}_{\theta}(P,W))^{\circ}$. Conversely, when $\tau^{\prime}$ in $\mathcal{C}_{\theta}(P,W)$ equals to 00, we can not get ${\rm trace}(\widetilde{H}_{\alpha\alpha})=0$, i.e., the reverse conclusion is not true. Then the proof of part (i) is completed.
@@
$\begin{split}\langle H,D\rangle&=\langle\widetilde{H},\widetilde{D}\rangle=\langle\widetilde{H}_{\alpha\alpha},\widetilde{D}_{\alpha\alpha}\rangle+\langle\widetilde{H}_{\alpha\beta},\widetilde{D}_{\alpha\beta}\rangle+\langle\widetilde{H}_{\beta\alpha},\widetilde{D}_{\beta\alpha}\rangle+\langle\widetilde{H}_{\alpha c},\widetilde{D}_{\alpha c}\rangle\\[10.0pt] &=\langle\widetilde{H}_{\alpha\alpha},T(\widetilde{D}_{\alpha\alpha})\rangle+\langle\widetilde{H}_{\alpha\beta},\widetilde{D}_{\alpha\beta}\rangle+\langle\widetilde{H}_{\beta\alpha},\widetilde{D}_{\beta\alpha}\rangle+\langle\widetilde{H}_{\alpha c},\widetilde{D}_{\alpha c}\rangle\\[10.0pt] &=\sum_{\begin{subarray}{c}1\leq l\leq r_{1}\\ 1\leq t\leq\widetilde{r}\end{subarray}}\frac{\nu_{l}(Q)-\lambda^{*}}{\lambda^{*}}\|(T(\widetilde{H}_{1}))_{a_{l}a_{t}}\|^{2}+\sum_{1\leq l\leq r_{1}}\frac{\nu_{l}(Q)-\lambda^{*}}{\lambda^{*}}\|\widetilde{H}_{a_{l}c}\|^{2}\\[4.0pt] &\quad\quad+\sum_{\begin{subarray}{c}1\leq l\leq r_{1}\\ \widetilde{r}+1\leq t\leq r+1\end{subarray}}\left(-\frac{2(\nu_{l}(Q)-\lambda^{*})}{\nu_{t}(Q)-\lambda^{*}}\|S(\widetilde{H}_{1})_{a_{l}a_{t}}\|^{2}+\frac{2(\nu_{l}(Q)-\lambda^{*})}{\nu_{t}(Q)+\lambda^{*}}\|T(\widetilde{H}_{1})_{a_{l}a_{t}}\|^{2}\right)\\[4.0pt] &=-\sum_{\begin{subarray}{c}1\leq l\leq r_{1}\\ \widetilde{r}+1\leq t\leq r+1\end{subarray}}\frac{2\sigma_{l}(W)}{\nu_{t}(P)-\lambda^{*}}\|S(\widetilde{H}_{1})_{a_{l}a_{t}}\|^{2}+\sum_{1\leq l\leq r_{1}}\frac{\sigma_{l}(W)}{\lambda^{*}}\|\widetilde{H}_{a_{l}c}\|^{2}\\[4.0pt] &\quad\quad+\sum_{\begin{subarray}{c}1\leq l\leq r_{1}\\ 1\leq t\leq r+1\end{subarray}}\frac{2\sigma_{l}(W)}{\nu_{t}(P)+\lambda^{*}}\|(T(\widetilde{H}_{1})_{a_{l}a_{t}}\|^{2},\\[4.0pt] \end{split}$
@@ 4 The Robust ICKKTM
In this section, owing to the special linear structure of the problem (1.1), we first show the primal SRCQ is equivalent to the dual SOSC. In fact, when the spectral norm in problem (1.1) is substituted by the nuclear norm, the same results are already given in [16]. Therefore, we only briefly state here, the main tools we used are proposition 3.1 and Proposition 3.2. The Lagrangian dual of problem (1.1) is as follows:
@@
Thus, we have that $(H_{1},0,0)\in\mathcal{C}(\bar{y},\bar{w},\overline{S})\backslash\{0\}$ and $\langle 0,(\nabla h^{*})^{\prime}(\bar{w};0)\rangle-\varphi^{*}_{(\overline{S},0)}(\overline{X})=0$, which contradicts the SOSC (4.5) at $(\bar{y},\bar{w},\overline{S})$ with respect to $\overline{X}$.
“(ii)$\Longrightarrow$(i)” Suppose on the contrary that the SOSC (4.5) does not hold at $(\bar{y},\bar{w},\overline{S})$ for $\overline{X}$. Then, $\exists\overline{H}=(\overline{H}_{y},\overline{H}_{w},\overline{H}_{S})\in\mathcal{C}(\bar{y},\bar{w},\overline{S})\backslash\{0\}$ such that
@@
From [25, Proposition 16] we know that $\varphi^{*}_{(\overline{S},H)}(\overline{X})\leq 0$ for any $H\in\Re^{m\times n}$. Moreover, because $h$ is essentially strictly convex, by [20, Theorem 26.3], $h^{*}$ is essentially smooth. Then, $\nabla h^{*}$ is locally Lipschitz continuous and directionally differentiable on ${\rm int}({\rm dom}\,h^{*})$. Since $(\bar{y},\bar{w},\overline{S})$ is the optimal solution, it is not hard to verify by the mean value theorem that $\langle H_{w},(\nabla h^{*})^{\prime}(\bar{w};H_{w})\rangle>0$ for any $H_{w}\neq 0$ such that $(H_{y},H_{w},H_{S})\in\mathcal{C}(\bar{y},\bar{w},\overline{S})\backslash\{0\}$. Therefore,
@@
From $\overline{H}\in\mathcal{C}(\bar{y},\bar{w},\overline{S})\backslash\{0\}$, we know that
@@
Switching the “primal” and “dual” in the above Theorem, owing to the special expression of $\mathcal{C}_{\theta}(P,W)$, we can only obtain that the primal SOSC is the sufficient condition for the Dual SRCQ from Proposition 3.2. One may refer to [16, Theorem 5.1] for more details.
@@
Let $\overline{X}\in\Re^{m\times n}$ be an optimal solution of problem (1.1) with $\mathcal{M}_{P}(\overline{X})\neq\emptyset$. Let $(\bar{y},\overline{S})\in\mathcal{M}_{P}(\overline{X})$. If the dual SRCQ holds at $\bar{y}$ for $\overline{X}$, i.e.,
@@
Then the primal SOSC holds at $\overline{X}$ for $(\bar{y},\overline{S})$, i.e.,
$\langle\mathcal{Q}D,\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}D\rangle-\psi^{*}_{(\overline{X},D)}(\overline{S})>0,\quad\forall D\in\mathcal{C}(\overline{X})\backslash\{0\},$ (4.8)
@@
In the following, we first give the definition of the KKT mapping for problem (1.1). Then, we demonstrate that the SOSC and the SRCQ are both sufficient and necessary for the ICKKTM. Finally, based on the above result, we give a series of characterizations of the robust ICKKTM for problem (1.1).
@@
$F(X,y,S):=\left[\begin{array}[]{c}C+\mathcal{B}^{*}y+\mathcal{Q}^{*}\nabla h(\mathcal{Q}X)+S\\[5.0pt] \mathcal{B}X-b-\Pi_{Q}(\mathcal{B}X-b+y)\\[5.0pt] X-{\rm Prox}_{\theta}(X+S)\end{array}\right],\quad(X,y,S)\in\Re^{m\times n}\times\Re^{l}\times\Re^{m\times n}.$ (4.9)
@@
$S_{\rm KKT}(\delta):=\{(X,y,S)\in\Re^{m\times n}\times\Re^{l}\times\Re^{m\times n}\mid F(X,y,S)=\delta\}.$ (4.10)
@@
$\langle\mathcal{Q}D,\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}D\rangle-\psi^{*}_{(\overline{X},D)}(\overline{S})>0,\quad\forall D\in\mathcal{C}^{0}(\overline{X})\backslash\{0\},$ (4.12)
@@
“(i)$\Longrightarrow$ (ii)”. Let $(\Delta Z,\Delta y,\Delta S)\in\Re^{m\times n}\times\Re^{l}\times\Re^{m\times n}$ satisfies
@@
$\Delta Z\in\mathcal{C}^{0}(\overline{X}).$ (4.16)
@@
$\begin{split}&\langle\Delta Z,\mathcal{Q}^{*}\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}\Delta Z+\mathcal{B}^{*}\Delta y+\Delta S\rangle\\[5.0pt] =&\langle\Delta Z,\mathcal{Q}^{*}\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}\Delta Z\rangle+\langle\Delta Z,\mathcal{B}^{*}\Delta y\rangle+\langle\Delta Z,\Delta S\rangle\\[5.0pt] =&\langle\mathcal{Q}\Delta Z,\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}\Delta Z\rangle+\langle\mathcal{B}\Delta Z,\Delta y\rangle+\langle\Delta Z,\Delta S\rangle\\[5.0pt] =&\langle\mathcal{Q}\Delta Z,\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}\Delta Z\rangle-\psi^{*}_{(\overline{X},\Delta Z)}(\overline{S})\\[5.0pt] =&0.\end{split}$ (4.17)
@@
We prove “(ii)$\Longrightarrow$ (i)” by contradiction. Assume that the SSRCQ (4.11) does not hold at $\overline{X}$ for $(\bar{y},\overline{S})$. Then $\exists(\Delta y,\Delta S)\neq 0$ s.t. (4.20) holds, which also means that (4.19) holds. Hence, $F^{\prime}((\overline{X},\bar{y},\overline{S});(0,\Delta y,\Delta S))=0$. This contradicts with the ICKKTM at the origin for $(\overline{X},\bar{y},\overline{S})$.
@@
$\langle\mathcal{Q}D,\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}D\rangle-\psi^{*}_{(\overline{X},D)}(\overline{S})\geq 0,\quad\forall D\in\mathcal{C}^{0}(\overline{X}).$ (4.22)
@@
$\langle\mathcal{Q}\Delta Z^{\prime},\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}\Delta Z^{\prime}\rangle-\psi^{*}_{(\overline{X},\Delta Z^{\prime})}(\overline{S})=0.$ (4.23)
@@
$\mathcal{Q}\Delta Z^{\prime}=0,\quad\psi^{*}_{(\overline{X},\Delta Z^{\prime})}(\overline{S})=0.$ (4.24)
@@
$F^{\prime}((\overline{X},\bar{y},\overline{S});(\Delta Z^{\prime},0,0))=0.$ (4.28)
@@
Finally, we can establish a series of sufficient and equivalent conditions of the robust ICKKTM for problem (1.1).
@@
Let $\overline{X}\in\Re^{m\times n}$and $(\bar{y},\bar{w},\overline{S})\in\Re^{l}\times\Re^{d}\times\Re^{m\times n}$ be optimal solutions of problem (1.1) and (4.1), respectively. Assume that $(\bar{y},\overline{S})\in\mathcal{M}_{P}(\overline{X})$ and $\overline{X}\in\mathcal{M}_{D}(\bar{y},\bar{w},\overline{S})$, and the RCQ (4.29) holds at $\overline{X}$. Consider the following statements:
@@
    The primal SOSC (4.8) holds at $\overline{X}$ for $(\bar{y},\overline{S})$ and the primal SSRCQ (4.11) holds at $\overline{X}$ for $(\bar{y},\overline{S})$.
@@
    The dual SRCQ (4.7) holds at $(\bar{y},\bar{w},\overline{S})$ for $\overline{X}$ and the primal SSRCQ (4.11) holds at $\overline{X}$ for $(\bar{y},\overline{S})$.
  • (g)
    The dual SOSC (4.5) holds at $(\bar{y},\bar{w},\overline{S})$ for $\overline{X}$ and the primal WSOSC (4.12) holds at $\overline{X}$ for $(\bar{y},\overline{S})$.
  • (h)
    The primal SRCQ (4.6) holds at $\overline{X}$ for $(\bar{y},\overline{S})$ and the primal WSOSC (4.12) holds at $\overline{X}$ for $(\bar{y},\overline{S})$.
  • (i)
  • (j)
@@
It holds that $(g)\Leftrightarrow(h)\Leftarrow(a)\Leftrightarrow(b)\Leftrightarrow(c)\Leftrightarrow(d)\Leftarrow(e)\Leftarrow(f)$, and $(i)\Leftrightarrow(j)\Rightarrow(a)$.
@@
Since the spectral norm function $\|\cdot\|_{2}$ is $\mathcal{C}^{2}$-cone reducible, the “robustly” in (i) of the Theorem 4.4, which means that $S_{\rm KKT}$ is also lower semicontinuous at the origin for $(\overline{X},\bar{y},\overline{S})$, can be obtained automatically from [13, Theorem 17]. Besides, since the spectral norm function is Lipschitz continuous at $\overline{X}$, by combining Theorem 4.1, Theorem 4.2 and [16, Proposition 3.3], we can obtain that $(a)\Leftrightarrow(c)\Leftrightarrow(d)\Leftarrow(i)\Leftrightarrow(j)$. Further, by combining Theorem 4.1, Theorem 4.2, Theorem 4.3 and [13, Theorem 17], we can derive that the remaining relationships hold true. ∎

### v3

@@ Abstract
This paper aims to provide a series of characterizations of the robust isolated calmness of the Karush-Kuhn-Tucker (KKT) mapping for spectral norm regularized convex optimization problems. By establishing the variational properties of the spectral norm function, we directly prove that the KKT mapping is isolated calm if and only if the strict Robinson constraint qualification (SRCQ) and the second-order sufficient condition (SOSC) hold. Furthermore, we obtain the crucial result that the SRCQ for the primal/dual problem and the SOSC for the dual/primal problem are equivalent. The obtained results can derive more equivalent conditions of the robust isolated calmness of the KKT mapping, thereby enriching the stability theory of spectral norm regularized optimization problems and enhancing the usability of isolated calmness in algorithm applications.
Keywords: Isolated calmness, Second-order sufficient condition, Spectral norm, Strict Robinson constraint qualification, Critical cone
@@ 1 Introduction
Stability analysis theory is very crucial in studying the convergence of algorithms. As one of the important concepts in stability properties, isolate calmness (see Definition 2.1) may guarantee the linear convergence rate of some algorithms, such as the alternating direction method of multipliers [9] and the proximal augmented Lagrangian method [10]. There are many publications on studying the isolated calmness of the KKT mapping (ICKKTM) for optimization problems. For the nonlinear semidefinite programming, Zhang and Zhang [11] show that the ICKKTM can be derived from the SOSC and the SRCQ. Zhang et al. [12] demonstrate that the SOSC and the SRCQ are both sufficient and necessary for the ICKKTM for the nonlinear second order cone programming problem. It is important to note that Ding, Sun and Zhang [13] study the robust isolated calmness of a large class of cone programming problems. They prove that if and only if both the SOSC and the SRCQ hold, the KKT mapping is robustly isolated calm.
@@
Inspired by the work in [16], given the widespread application of spectral norm regularized convex matrix optimization problems, a natural question is whether the results in [16] can be extended to spectral norm regularized convex optimization problems. We provide an positive answer in this paper. Due to the special structure of the critical cone of spectral norm function, we provide several equivalent conditions of the robust ICKKTM. That is to say, the conclusions in [16] are still valid for problem (1.1). Compared to [16], the difference in this paper is that we use the established variational properties of the proximal mapping of spectral norm to provide a direct proof of the ICKKTM for problem (1.1), while [16] provides an indirect proof of the isolated calmness based on [13, Theorem 24] regarding the optimization problems with a $\mathcal{C}^{2}$-cone reducible constraint set. In addition, compared to the SOSC, the SRCQ for primal or dual problem is easier to verify, which can enhance the practicality of isolated calmness in algorithm research.
The organization of the subsequent content is as follows. We provide some notations and preliminaries on variational analysis which will be used in the following text in Section 2. The variational properties of the spectral norm function are studied in Section 3, including the relationship between the critical cones of the spectral norm function and its conjugate function, and the explicit expression of the directional derivative of the proximal mapping of the spectral norm function. The results in Section 3 play a vital role in obtaining the important conclusions of this paper. In Section 4, we prove that the primal(dual) SRCQ holds if and only if the dual(primal) SOSC holds and thus establish more equivalent conditions for the robust ICKKTM for problem (1.1). We conclusion this paper in Section 5.
Some common symbols and notations for matrices are as follows:
@@ 2 Notation and Preliminaries
Moreover, we say that $G$ is robustly isolated calm at $\bar{u}$ for $\bar{v}$ if (2.1) holds, and for each $u\in\mathcal{U}$, $G(u)\cap\mathcal{V}\neq\emptyset$.
For any closed convex function $g:\mathcal{X}\rightarrow(-\infty,+\infty]$, we know from [18, Proposition 2.58] that $g$ is directionally epidifferentiable. We use $g^{\downarrow}(x;\cdot)$ to denote the directional epiderivatives of $g$. Further, if $g^{\downarrow}(x;d)$ is finite for $x\in{\rm dom}\,g$ and $d\in\mathcal{X}$, we define the lower second-order directional epiderivative of $g$ for any $w\in\mathcal{X}$ by
@@ 3 Variational Analysis of the Spectral Norm Function
In this section, we will provide a direct expression for the directional derivative of the proximal mapping of the spectral norm function. Then discuss the relationship between directional derivatives of the proximal mapping, the so-called “sigma term” and the critical cones.
@@
where $U\in\mathcal{O}^{m}$ and $V_{1}\quad V_{2}]\in\mathcal{O}^{n}$ with $V_{1}\in\Re^{n\times m}$ and $V_{2}\in\Re^{n\times(n-m)}$. Define the following three index sets
@@
where $\lambda^{*}>0$ and satisfies $\sum\limits_{i=1}^{m}[\sigma_{i}(Q)-\lambda^{*}]_{+}=1$ (for any $\in\Re, [x]+[x$$_{+}$ means the nonnegative part of $x$). Then, we can conclude that $\kappa(\sigma)=(\phi(\sigma_{1}),\phi(\sigma_{2}),\cdots,\phi(\sigma_{m}))$ and
@@
$\alpha:=\bigcup_{k=1}^{\tilde{r}}a_{k},\quad\beta:=\bigcup_{k=\tilde{r}+1}^{r+1}a_{k}.$ (3.4)
@@
[…]
@@
Propositions 10 and 12 in [25] characterized the critical cones of $\theta$ and $\theta^{*}$ defined in (3.8) and (3.9), respectively. We summarize the results as follows.
@@
    where $\lambda(Z)$ denote the real eigenvalue vector of symmetric matrix $Z$ being arranged in nonincreasing order.
@@
    If $H\in\mathcal{C}_{\theta}(P,W)$ and $\psi^{*}_{(P,H)}(W)=0$ if and only if $H\in(\mathcal{C}_{\theta^{*}}(P,W))^{\circ}$.
@@
Then, it follows from part (ii) of Proposition 3.1 that $H\in\mathcal{C}_{\theta^{*}}(W,P)$ and $\varphi^{*}_{(W,H)}(P)=0$ are equivalent to
@@
which means that $H\in(\mathcal{C}_{\theta}(P,W))^{\circ}$. Conversely, when $\tau^{\prime}$ in $\mathcal{C}_{\theta}(P,W)$ equals to 0, we can not get ${\rm trace}(\widetilde{H}_{\alpha\alpha})=0$, i.e., the reverse conclusion is not true. Then the proof of part (i) is completed.
@@
[…]
@@ 4 The Robust ICKKTM
In this section, owing to the special linear structure of the problem (1.1), we first show the primal (dual) SRCQ is equivalent to the dual (primal) SOSC. In fact, when the spectral norm in problem (1.1) is substituted by the nuclear norm, the same results are already given in [16]. Therefore, we only briefly state here, and the main tools we used are propositions 3.1 and 3.2. The Lagrangian dual of problem (1.1) is as follows:
@@
[…]
@@
[…]
@@
[…]
@@
Switching the “primal” and “dual” in the above Theorem, owing to the special expression of $\mathcal{C}_{\theta}(P,W)$, we can also obtain that the primal SOSC is equivalent to the dual SRCQ from Proposition 3.2. One may refer to [16, Theorem 5.1] for more details.
@@
Let $\overline{X}\in\Re^{m\times n}$ be an optimal solution of problem (1.1) with $\mathcal{M}_{P}(\overline{X})\neq\emptyset$. Let $(\bar{y},\overline{S})\in\mathcal{M}_{P}(\overline{X})$. Then the following results are equivalent:
  • (i)
    The dual SRCQ holds at $\bar{y}$ for $\overline{X}$, i.e.,
@@
  • (ii)
    The primal SOSC holds at $\overline{X}$ for $(\bar{y},\overline{S})$, i.e.,
[…]
@@
In the following, we first give the definition of the KKT mapping for problem (1.1). Then, we demonstrate that the SOSC and the SRCQ are sufficient and necessary for the ICKKTM. Finally, based on the above result, we give a series of characterizations of the robust ICKKTM for problem (1.1).
@@
$F(X,y,S):=\left[\begin{array}[]{c}C+\mathcal{B}^{*}y+\mathcal{Q}^{*}\nabla h(\mathcal{Q}X)+S\\[5.0pt] \mathcal{B}X-b-\Pi_{Q}(\mathcal{B}X-b+y)\\[5.0pt] X-{\rm Prox}_{\theta}(X+S)\end{array}\right],\quad(X,y,S)\in\Re^{m\times n}\times\Re^{l}\times\Re^{m\times n}.$
@@
$S_{\rm KKT}(\delta):=\{(X,y,S)\in\Re^{m\times n}\times\Re^{l}\times\Re^{m\times n}\mid F(X,y,S)=\delta\}.$
@@
[…]
@@
[…]
@@
$\Delta Z\in\mathcal{C}^{0}(\overline{X}).$
@@
$\begin{split}&\langle\Delta Z,\mathcal{Q}^{*}\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}\Delta Z+\mathcal{B}^{*}\Delta y+\Delta S\rangle\\[5.0pt] =&\langle\Delta Z,\mathcal{Q}^{*}\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}\Delta Z\rangle+\langle\Delta Z,\mathcal{B}^{*}\Delta y\rangle+\langle\Delta Z,\Delta S\rangle\\[5.0pt] =&\langle\mathcal{Q}\Delta Z,\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}\Delta Z\rangle+\langle\mathcal{B}\Delta Z,\Delta y\rangle+\langle\Delta Z,\Delta S\rangle\\[5.0pt] =&\langle\mathcal{Q}\Delta Z,\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}\Delta Z\rangle-\psi^{*}_{(\overline{X},\Delta Z)}(\overline{S})\\[5.0pt] =&0.\end{split}$
@@
[…]
@@
$\langle\mathcal{Q}D,\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}D\rangle-\psi^{*}_{(\overline{X},D)}(\overline{S})\geq 0,\quad\forall D\in\mathcal{C}^{0}(\overline{X}).$
@@
$\langle\mathcal{Q}\Delta Z^{\prime},\nabla^{2}h(\mathcal{Q}\overline{X})\mathcal{Q}\Delta Z^{\prime}\rangle-\psi^{*}_{(\overline{X},\Delta Z^{\prime})}(\overline{S})=0.$
@@
$\mathcal{Q}\Delta Z^{\prime}=0,\quad\psi^{*}_{(\overline{X},\Delta Z^{\prime})}(\overline{S})=0.$
@@
$F^{\prime}((\overline{X},\bar{y},\overline{S});(\Delta Z^{\prime},0,0))=0.$
@@
Finally, we can establish a series of equivalent conditions of the robust ICKKTM for problem (1.1).
@@
Let $\overline{X}\in\Re^{m\times n}$and $(\bar{y},\bar{w},\overline{S})\in\Re^{l}\times\Re^{d}\times\Re^{m\times n}$ be optimal solutions of problem (1.1) and (4.1), respectively. Assume that $(\bar{y},\overline{S})\in\mathcal{M}_{P}(\overline{X})$ and $\overline{X}\in\mathcal{M}_{D}(\bar{y},\bar{w},\overline{S})$, and the RCQ (4.19) holds at $\overline{X}$. Then the following statements are equivalent:
@@
Since the spectral norm function $\|\cdot\|_{2}$ is $\mathcal{C}^{2}$-cone reducible, the “robustly” in (i) of the Theorem 4.4, which means that $S_{\rm KKT}$ is also lower semicontinuous at the origin for $(\overline{X},\bar{y},\overline{S})$, can be obtained automatically from [13, Theorem 17]. Besides, since the spectral norm function is Lipschitz continuous at $\overline{X}$, by combining Theorem 4.1, Theorem 4.2, Theorem 4.3, [13, Theorem 17] and [16, Proposition 3.3], we can obtain that $(a)\Leftrightarrow(b)\Leftrightarrow(c)\Leftrightarrow(d)\Leftrightarrow(e)\Leftrightarrow(f)$. ∎